### v1

# LA-MARRVEL: A Knowledge-Grounded and Language-Aware LLM Reranker for AI-MARRVEL in Rare Disease Diagnosis

Abstract
Diagnosing rare diseases often requires connecting variant-bearing genes to evidence that is written as unstructured clinical prose, which the current established pipelines still leave for clinicians to reconcile manually. To this end, we introduce LA-MARRVEL, a knowledge-grounded and language-aware reranking layer that operates on top of AI-MARRVEL: it supplies expert-engineered context, queries a large language model multiple times, and aggregates the resulting partial rankings with a ranked voting method to produce a stable, explainable gene ranking. Evaluated on three real-world cohorts (BG, DDD, UDN), LA-MARRVEL consistently improves Recall@K over AI-MARRVEL and established phenotype-driven tools such as Exomiser and LIRICAL, with especially large gains on cases where the first-stage ranker placed the causal gene lower. Each ranked gene is accompanied by LLM-generated reasoning that integrates phenotypic, inheritance, and variant-level evidence, thereby making the output more interpretable and facilitating clinical review.
Introduction
Rare diseases affect millions of people1, yet each single condition is uncommon and often hard to diagnose. Many patients wait 4.7 years for an answer2. Modern DNA and RNA tests can identify genes that may explain a person’s symptoms, but they typically return a long list of candidates 3, 4. Clinicians then have to verify across papers, databases, and case reports to decide which genes truly match the patient. This step is labor-intensive and time-consuming, and it is easy to overlook critical details embedded within plain language5.
Large language models (LLMs) can read and explain complex medical text. For example, the recent models show higher than 95% of accuracy for a USMLE (United States Medical Licensing Examination)-style benchmark 6, 7. However, we previously showed that LLMs may exhibit bias toward well-studied genes and sensitivity to the ordering of candidate genes, which can compromise the reliability of gene-ranking results 8. These limitations suggest that, for clinical decision-support tasks, LLMs requires grounding and stability: linking outputs to citable knowledge (grounding) and producing consistent responses across repeated runs (stability). Without these properties, even strong language skills are not enough for safe clinical use 9, 10. A major reason for this challenge is that medical knowledge is largely written as free text; for example, about 80% of all clinical data are found in this unstructured form 11. Even when symptoms and inheritance are encoded using ontologies, converting them into simple tabular formats for traditional machine learning still demands extensive normalization and feature engineering12 or loses important context. The computational gene-prioritization tools for rare-disease diagnosis are typically used as a high-recall first stage to filter/rank an initial list of candidate genes, identifying many plausible targets but still leaving a few false positives
citepJacobsen2022. Turning that a long gene list into a short, trustworthy set of likely genes requires careful reading of unstructured text and cross-checking sources —the kind of labor-intensive task that diverts valuable time and attention from busy clinical practices13, 14.
To bridge these gaps, we introduce LA-MARRVEL, a language-aware reranking system that sits on top of a knowledge-driven high-recall pipeline. Rather than supplanting established bioinformatics tools, our goal is to further prioritize candidate genes, yielding a more concise, accurate, and defensible list. LA-MARRVEL brings together three pieces. First, we provide expert-engineered context: curated phenotype and disease information that gives the model the right facts at the right time. Second, we use a ranked voting method (Tideman’s method)15 to combine multiple responses into a single, consensus gene order. The ranked voting method decreases randomness and favors choices supported by repeated evidence. Lastly, we integrate this layer with the AI-MARRVEL pipeline 16, which already produces strong first-stage ranks and gene annotations. (Figure˜1)
We evaluate LA-MARRVEL on three real-world patient cohorts —Baylor Genetics (BG), Deciphering Developmental Disorders (DDD), and the Undiagnosed Diseases Network (UDN). These datasets come from independent cohorts which are representative of cases commonly encountered in clinical practice. We compare our approach to widely used diagnostic tools (e.g., Exomiser17 and LIRICAL18) and to general-purpose LLMs. Across the three cohorts, we demonstrated LA-MARRVEL improves top-rank accuracy without compromising recall, resulting in more reliable prioritization of the causal gene and consequently reducing the downstream review burden.
A notable advantage of LA-MARRVEL is its transparency. For each ranked gene, the system provides plain-language explanations of how the patient’s phenotypes correspond to the gene’s known disease features and how the observed variants accord with the expected mode of inheritance. As the explanation is derived from expert knowledge and produced using a consensus-based methodology, LA-MARRVEL helps clinicians can more confidently evaluate and rely on the analytic result. Consequently, clinician effort can be redirected from preliminary data triage toward substantive clinical decision.
Results
LA-MARRVEL outperforms gene-prioritization algorithms across three independent datasets
We benchmarked LA-MARRVEL against widely used phenotype-driven gene-prioritization amethods (Exomiser and LIRICAL) and an internal baseline (AI-MARRVEL) on three independent cohorts —BG, DDD, and UDN —using Recall@K (the fraction of cases in which the causal gene appears within the top-K ranked candidates). (Figure˜2)
Across all datasets and all K values (K=1-10), LA-MARRVEL consistently achieved the highest recall. At the clinically salient low-K thresholds, LA-MARRVEL showed the largest gains: at Top-1 and Top-3 it improved by roughly 5-10 percentage points over AI-MARRVEL, 20-30 points over Exomiser, and 30-45 points over LIRICAL. The advantage is especially pronounced on UDN, where LA-MARRVEL reaches ~90% by Top-3 while the next best method trails several points lower. On BG and DDD, LA-MARRVEL approaches a ceiling rapidly, surpassing ~95% by Top-5 and maintaining a small but consistent margin over AI-MARRVEL at every K.
Performance at larger K also favors LA-MARRVEL. By Top-10, it attains ~95-98% recall on all three cohorts, compared with ~85-90% for Exomiser and ~75-80% for LIRICAL. These results indicate that LA-MARRVEL not only surfaces the correct diagnosis more often at the very top of the list—where it matters most for clinician time and follow-up testing —but also maintains superior coverage when a wider gene set is considered.
Together, these findings demonstrate that LA-MARRVEL reliably generalizes across diverse cohorts and outperforms established gene-prioritization algorithms, offering both higher precision at the top ranks and near-ceiling recall with modest review effort.
Larger Performance Gains Among Initially Poorly-Ranked Genes
Across all of the three cohorts (BG, DDD, UDN), the LA-MARRVEL shows its largest absolute gains for examples where the original ranker placed the target further down the list (i.e., poor rank). When the baseline already puts the right item at or near the top, there is less room to help, and occasional harms appear, but when the baseline misses the top positions, the re-ranker very often rescues the example. (Figure˜3)
Key observations are:
  • Strong rescue at poorer ranks. For BG the bars for original ranks 3-9 are entirely green (100% improved; small $n$), and for DDD and UDN most ranks $\geq 3$ are also 100% improved. That indicates LA-MARRVEL is very effective at promoting correct candidates that the baseline ranked lower.
  • Some harm concentrated at top positions. At original rank 0 the harmed fraction is non-negligible (BG ~17.1%, DDD ~10.4%, UDN ~10.1%). Ranks 1-2 also show the largest harmed/neutral shares compared with deeper positions. This suggests the re-ranker sometimes demotes an already-correct top result when the baseline was already good.
In summary, LA-MARRVEL delivers its largest improvements when the original ranker places the correct item further down the list.
LLMs Require a First-Stage Ranker for Effective Prioritization.
Directly prompting state-of-the-art LLMs to name causal genes performs poorly compared with rank-then-rerank pipelines. In our evaluation (Figure˜4), prompt-only LLMs (Claude and Claude-Thinking, both without AIM) recover only ~12-15% of cases at Recall@1, far below Exomiser (~50%), LIRICAL (~31%), and LA-MARRVEL (~78%). The gap persists at wider review windows: by Top-5/Top-10, LLM-only recall remains ~37-55% versus ~75-85% for classical tools and ~95-98% for LA-MARRVEL; even at Top-100 they plateau around ~70-76% compared with ~96-99% for the other methods.
These results indicate that LLMs are not reliable stand-alone retrieval engines for rare-disease gene discovery, when they are given the exhaustive list of all variant-bearing genes for each patient (benign and pathogenic alike). Faced with a long-tail of gene–phenotype links and a noisy, unfiltered candidate set, they miss true genes at clinically useful top-k cutoffs; free-form generation adds off-list distractors and hallucinated rationales; and reasoning alone ("Thinking") yields only modest gains. By contrast, pairing an LLM with a high-recall first-stage ranker (AI-MARRVEL) focuses the model on a comprehensive, evidence-grounded candidate set, enabling phenotype-aware judgment where it excels. Consequently, LA-MARRVEL’s rank-then-rerank design is essential for achieving high precision at the top ranks while maintaining near-ceiling recall.
Knowledge-Grounded Prompt Generation Enables Accurate LA-MARRVEL.
We observed that removing phenotype knowledge (Human Phenotype Ontology, HPO)19 from prompts causes the largest drop in retrieval accuracy, especially at the clinically crucial top ranks. At Top-1, the Recall@1 loss for Without HPO Information is 20.22%p, far exceeding Without Disease Information (5.90%p), Without Disease Phenotypes But with Name (4.78%p), and Without HPO Text Description But with ID (3.37%p). The pattern holds across $K$: at Recall@3, No HPO loses 10.39%p versus 1.97%p, 1.12%p, and 0.56%p for the other ablations, respectively. In short, removing phenotype knowledge produces the single largest degradation in performance. (Figure˜5)
Practical Implications are:
  • Top-rank sensitivity: Losses are greatest at K=1-3 and taper with larger K, so errors from missing knowledge are most damaging where precision matters most.
  • HPO is critical: Omitting HPO entirely is the worst mistake; other omissions hurt less.
  • Disease name alone is insufficient: Supplying only a disease name yields only a modest improvement over omitting disease information altogether. The reason is that many disease names such as Bohring-Opitz syndrome are not self-explanatory of symptoms.
  • HPO IDs carry most of the value: When HPO IDs are retained but their human-readable text is removed (No HPO Text), losses are small (e.g., 3.37%p at K=1), indicating the HPO ID itself is highly informative.
In summary, LA-MARRVEL’s accuracy depends on knowledge-grounded prompt generation. Incorporating phenotype signals (HPO), linking explicit disease context to well-specified phenotypes, and engineering context to highlight disambiguating details are the most effective ways to reduce recall loss at clinically meaningful top ranks.
The Number of Considered Genes has Trade-Offs
Increasing the number of candidate genes ($G$) shifts retrieval behavior in two opposing ways. With fewer candidates, the model is more decisive and more often places the correct gene at rank 1 boosting Recall@1. As $G$ grows, the pool broadens and the system captures more true positives within the top 10 raising Recall@10. In our experiments, ($G$ = 20, 30, 50, 100) consistently improved Recall@10 over the first-stage ranker (AI-MARRVEL) simply because more plausible genes were evaluated. (Figure˜6)
This gain comes with a trade-off. When the baseline already ranks the correct gene near the top, adding many extra candidates can perturb the ordering and occasionally push the ground-truth gene lower. Performance improves most on harder cases (where the original rank was worse), while some easier cases show slight declines as $G$ increases. Fitted trends against the original first-stage ranks have positive slopes ($\approx$ 0.78 for $G=10$; $\approx$ 0.62 for $G=50$), indicating that the benefits grow with difficulty—larger $G$ helps recover missed genes but can slightly reduce performance on a few initially correct top-1 predictions.
In short, more candidates increase coverage but can dilute rank concentration. LA-MARRVEL pipeline exposes $G$ as an optional parameter, allowing users to tune for their goal: keep $G$ small to "find the single best gene," or increase $G$ to "cover the likely set within the top-10."
Ranked Voting consistently outperforms Single Run, with gains increasing in larger gene-sets
Across all gene-set sizes ($G$ = 10, 20, 30, 50, 100) and for every K, Ranked Voting (10x) yields higher Recall@K than a Single Run (1x). The advantage is visible from small K (Top-2/Top-3) through larger K, and the curves do not cross —Ranked Voting — i.e., always perform better than Single Run. (Figure˜7)
The magnitude of this gain grows with $G$. With few genes (e.g., GG=10), the curves are close, but by (GG=50,100) the separation is pronounced even at modest K, and Single Run can underperform the first-stage ranker. This pattern is consistent with our expectations: larger candidate sets introduce higher per-run variance; voting aggregates signal across runs, dampening noise and surfacing true positives more reliably. Therefore, the benefit of ensembling scales with problem size.
LA-MARRVEL provides an Explainable Trace
To facilitate the review of gene rankings, LA-MARRVEL generates an explainer trace — a step-by-step view that adds evidence to each candidate gene. Each panel shows ACMG evidence codes — the American College of Medical Genetics and Genomics criteria for classifying pathogenic variants20 — zygosity checks, and how well the phenotype matches HPO terms. This trace follows the manual curation process and records the reasons for any changes in rank.
Case Study 1: SPG7 reprioritized to highest rank
In this case, AI-MARRVEL put CLDN16 at rank 1 and SPG7 at rank 2, even though the true diagnosis is linked to SPG7 (Figure˜8). LA-MARRVEL’s reasoning trace walked through the same evidence a human would check and corrected the order.
LLM Explainer Trace for Case Study 1 […] CLDN16 ACMG Classification: Likely Benign Molecular evidence: Very low CADD (0.019) suggests benign Single heterozygous variant for AR condition ClinVar: Uncertain significance Phenotype matching: POOR (3/10) Primarily renal/metabolic disorder Some overlap: nystagmus, muscle weakness, hypertension Conclusion — Insufficient evidence for pathogenicity SPG7 ACMG Classification: Likely Pathogenic Molecular evidence: High impact missense (CADD 27.8 — strong computational evidence) ClinVar: Conflicting interpretations suggests some pathogenic evidence Concern: High population frequency (0.002923) reduces pathogenicity likelihood Phenotype matching: EXCELLENT (9/10) HPO matches: Nystagmus, Muscle weakness, Gait ataxia, Dysarthria, Dysdiadochokinesis, Spasticity, Unsteady gait Disease inheritance: AD/AR (heterozygous variant can be causative in AD form) Conclusion — Despite population frequency concerns, phenotype match is exceptional [additional genes omitted for brevity]
Why CLDN16 was demoted.
The reasoning trace identified multiple mismatches for CLDN16. The variant itself looks weak: it has a very low CADD score (0.019), is classified in ClinVar as of uncertain significance, and is present in only a single heterozygous state even though the associated condition is autosomal recessive. On top of that, the patient’s phenotype does not capture the core renal/metabolic manifestations of CLDN16-related disease (hypomagnesemia, polyuria, hypercalciuria, CKD); the overlap is limited to a few nonspecific findings like nystagmus or muscle weakness. Given weak molecular support, inappropriate zygosity for an AR disorder, and poor overall phenotype fit, LA-MARRVEL pushed CLDN16 down the list.
Why SPG7 was promoted.
In contrast, SPG7 showed a pattern that supports causality. The variant has strong computational support (missense with CADD 27.8), and although ClinVar reports conflicting interpretations, the trace treated this as compatible with some pathogenic evidence. Most importantly, the phenotype match was rated excellent (9/10): spasticity, gait ataxia, unsteady gait, dysarthria, dysdiadochokinesis, nystagmus, and muscle weakness all map cleanly to spastic paraplegia 7. Because SPG7 is described in both AD and AR contexts, the presence of a single heterozygous variant was not taken as disqualifying. With much stronger phenotype alignment and plausible pathogenicity despite a population-frequency concern, LA–MARRVEL reordered the candidates and placed SPG7 first.
Case Study 2: STIM1 reprioritized to highest rank
In this case, AI-MARRVEL initially ranked UNC80 first and STIM1 ninth, even though the ground truth is linked to STIM1 (Figure˜9). LA-MARRVEL’s explainable trace re-evaluated the molecular, inheritance, and phenotype evidence and corrected the ranking. “‘
LLM Explainer Trace for Case Study 2 […] UNC80 ACMG Assessment Uncertain Significance Variant Quality: Very high CADD (30.0), very rare AF (0.0), but ClinVar uncertain significance Inheritance Mismatch: Autosomal recessive condition but only heterozygous variant found Phenotype Matching (Excellent but inheritance doesn’t fit): Nystagmus, feeding difficulties, absent speech, hypoplasia of corpus callosum, global developmental delay, short philtrum, esotropia, seizure, prominent nasal bridge, intrauterine growth retardation Conclusion: Strong phenotype match but inheritance pattern incompatible [genes omitted for brevity] STIM1 ACMG Assessment: Likely Pathogenic Variant Quality: High CADD (24.9), rare AF (4.375e-05), missense with MODERATE impact Inheritance Match: Heterozygous variant compatible with autosomal dominant Stormorken syndrome Phenotype Matching (Excellent): Epistaxis (nosebleeds) Thrombocytopenia Elevated circulating creatine kinase concentration Short stature Deeply set eye Short philtrum Conclusion: Strong candidate — inheritance pattern fits, multiple phenotype matches [additional genes omitted for brevity]
Why UNC80 was demoted.
The initial AI-MARRVEL output favored UNC80 because many of the patient’s HPO terms lined up with reported UNC80-related neurodevelopmental disease (nystagmus, feeding difficulty, developmental delay, corpus callosum hypoplasia). However, LA-MARRVEL’s trace treated this as incomplete because the molecular and inheritance context did not support causality. The disorder associated with UNC80 is classically autosomal recessive, but only a single heterozygous variant was present. Even though the variant itself looked compelling on paper (CADD 30.0, absent from population databases), the lack of a second hit means the genotype does not meet the expected disease model. ClinVar also did not provide strong confirmatory evidence (uncertain significance), so there was no external support to override the inheritance mismatch. LA-MARRVEL therefore down-weighted UNC80: strong phenotype match alone was not enough to keep it at rank 1 when the zygosity and inheritance pattern were incompatible.
Why STIM1 was promoted.
In contrast, STIM1 satisfied all three axes that LA-MARRVEL tracks: variant plausibility, inheritance, and phenotype fit. The variant was rare (AF $4.375\times 10^{-5}$), missense, and computationally supported (CADD 24.9), which is consistent with a pathogenic or likely pathogenic call in an autosomal dominant context. The trace explicitly noted that a single heterozygous STIM1 variant is compatible with Stormorken syndrome, so unlike UNC80, the genotype matched the known disease mechanism. Most importantly, the phenotype alignment was unusually specific: epistaxis, thrombocytopenia, elevated CK, short stature, and characteristic facial features (deep-set eyes, short philtrum) all map to the STIM1-related phenotypes. Because LA-MARRVEL prioritizes candidates where all lines of evidence agree, it upgraded STIM1 from rank 9 to the top position: rare, plausible variant; correct AD inheritance; and a multi-system phenotype that is far more characteristic than the broader neurodevelopmental presentation that pushed UNC80 up initially.
Discussion
LA-MARRVEL is built on three key ideas: i) reranking with an LLM on top of a high-recall bioinformatics pipeline is a better fit for rare-disease diagnosis than using an LLM alone, ii) language-aware, knowledge-grounded prompts are necessary because much clinical signal lives in prose (HPO descriptions, inheritance notes, atypical presentations), and iii) stability can be achieved by aggregating multiple LLM judgments through a deterministic voting procedure. With this design, LA-MARRVEL produces shorter and better-justified gene lists than the first-stage system, preserves near-ceiling recall, and exposes an explanation trace that matches how clinicians already review cases. Empirically, this lets us outperform established diagnostic tools such as AI-MARRVEL, Exomiser, and LIRICAL at clinically salient top-k cutoffs while still integrating cleanly into existing rare disease workflows.
While we took a first step toward language-aware reranking, several aspects motivate future work. First, our results make clear that knowledge-grounded prompt composition— especially inclusion of HPO IDs and disease-specific phenotype context—is the single largest driver of accuracy; removing phenotype knowledge produced the sharpest drop in Recall@1. This suggests LA-MARRVEL can be extended to richer prompt builders that pull from additional structured and semi-structured sources (clinic notes, longitudinal phenotype updates, or local variant databases) to mitigate missing or noisy HPO annotations. Second, we showed that ranked voting (Tideman) reduces per-run variance and becomes more valuable as the candidate gene set grows, but this comes at extra inference cost; sampling-efficient or adaptive ensembling strategies are a natural next step to keep the system responsive in production settings. Third, although our explainer trace already records phenotype matching, zygosity checks, and ACMG-like evidence, deeper clinical explainability—e.g., surfacing the exact citable source used in the prompt or aligning to institution-specific reporting templates—would make the system more directly reviewable by clinicians.
As the rare-disease stack evolves from single-model tools to compound systems that combine variant annotation, phenotype normalization, retrieval, and LLM reasoning, we need methods that “snap in” to existing bioinformatics pipelines rather than replace them. LA-MARRVEL fits this trajectory: it treats AI-MARRVEL and other first-stage rankers as authoritative for recall, then applies language-aware prioritization to win precision at the top of the list. Going forward, we see several concrete extensions: incorporating tool-use or retrieval-augmented steps to auto-cite knowledge sources, supporting multi-gene/oligogenic diagnoses that our current evaluation filtered out, and exploring meta-optimization of the reranking procedure itself (e.g., tuning prompt templates or voting parameters on held-out clinical cohorts). Finally, like other LLM-powered clinical systems, the ultimate test of LA-MARRVEL lies in prospective, real- world clinical assessments and in integration with institutional data and review processes, which are outside the scope of this paper.
Materials and Methods
Datasets
We reuse three clinical cohorts from AI-MARRVEL to ensure a consistent, apples-to-apples comparison with prior work. Summary statistics are shown in Table˜1. In brief:
  • BG (Baylor Genetics). 63 cases; on average 1469.2 candidate genes, 10.59 HPO terms, and 1.175 causal genes per case, spanning 74 unique causal genes. Curated in-house; accessibility is Internal.
  • DDD (Deciphering Developmental Disorders). 214 cases; on average 811.5 candidate genes, 7.38 HPO terms, and 1.000 causal genes per case, spanning 116 unique causal genes. Curated by AMELIE; accessibility is Restricted.
  • UDN (Undiagnosed Diseases Network). 90 cases; on average 1072.2 candidate genes, 43.98 HPO terms, and 1.044 causal genes per case, spanning 93 unique causal genes. Curated in-house; accessibility is Restricted.
We adopt the original curation provided by each source, but we further filtered out multi-gene cases to provide error analysis.
Experiment Details
Model For all LA-MARRVEL experiments, we used Claude-4-Sonnet with Extended Thinking enabled (anthropic.claude-sonnet-4-20250514-v1:0). The model was configured with a 65,536-token context window and a 50,000-token thinking budget for each call.
Hyperparameters Unless otherwise specified, all decoding and sampling hyperparameters (e.g., temperature, top_k, top_p, and related settings) were kept at their default values provided by the Bedrock-hosted Claude-4-Sonnet endpoint. No additional tuning of these hyperparameters was performed across experiments.
Environment All experiments were conducted in a HIPAA-compliant Amazon Web Services (AWS) environment, accessing the model exclusively via the Amazon Bedrock API. During evaluation, the model was run in a strictly offline configuration with respect to external knowledge sources: no web search, browsing, or retrieval from external APIs was enabled or used in any of the LA-MARRVEL experiments.
First-Stage Ranks and Gene Annotations by AI-MARRVEL
LLMs are evolving, but at the current stage they require concise, informative context provided via a prompt. On average, a proband has variants in 800-1,500 genes, including benign ones; including all genes in a prompt is not feasible. To reduce the number of candidate genes to a feasible size, we chose the reputable tool AI-MARRVEL, known for its state-of-the-art methods in rare disease diagnosis on the BG, DDD, and UDN cohorts 16.
Knowledge-Grounded Prompt Composer
AI-MARRVEL not only ranks candidate genes, it also generates variant- and gene-level annotations using Ensembl’s Variant Effect Predictor (VEP) and the MARRVEL database. LA-MARRVEL then uses these to create concise gene summaries (See the below prompt template, Table˜2, and Table˜3), which can help preliminarily assess several ACMG/AMP criteria 20, including: variant consequence (PVS1, when predicted loss-of-function is a known disease mechanism); computational predictions such as CADD (PP3/BP4); gnomAD allele frequency (PM2/BA1, absent or rare / too common); gene-disease information (e.g., OMIM) combined with a highly specific patient phenotype (PP4); ClinVar (PP5/BP6, reputable database).
LA-MARRVEL Main Prompt Template Given HPO and Genes with Variants, Evaluate Gene 1 by 1 based on ACMG and phenotype matching to Prioritize Pathogenic Genes. Notes: zyg=het & isTransHeterozygote=yes: a second pathogenic variant is present in both alleles of a gene, which meets the recessive disease criteria. zyg=hom may mean hemizygous for chromosome X if the proband is male The data is generated from whole exome sequencing, finding additional pathogenic variant is unlikely. HPO: <HPO table> Genes: <Gene summary table>
Aggregating Partial Gene Rankings with Tideman’s Method
Ensembling LLMs at inference time —via self-consistency (sample diverse chains of thought and pick the modal answer), structured search like Tree-of-Thoughts, or Best-of-$N$ selection with a verifier —consistently boosts accuracy on math, code, and reasoning benchmarks 21, 22, 23. In addition, sampling-efficient variants such as ConSol use Sequential Probability Ratio Testing (SPRT) to adaptively stop once responses converge, cutting the number of samples required relative to vanilla self-consistency while preserving accuracy 24.
Here, each sample yields a partial ranking of candidate genes (unranked items are treated as missing). We aggregate these ballots using Ranked Pairs (Tideman) 15: (1) build pairwise win counts $\text{Wins}[a,b]$ by treating any ranked item as preferred over any candidate genes; (2) for each unordered pair, form a directed victory with margin $|\text{Wins}[a,b]-\text{Wins}[b,a]|$; (3) sort directed pairs by decreasing margin, then by the winner’s votes, with deterministic lexicographic tie-breaks; (4) greedily lock edges that do not create cycles; (5) output a topological order, breaking ties by Borda score (computed only from ranked positions) then by name (see Algorithm˜1 and 2). Ranked Pairs is Condorcet-compliant and clone-independent, leveraging full list structure while remaining robust to near-duplicate options.
Acknowledgments
This work was supported by the Cancer Prevention and Research Institute of Texas (CPRIT, RP240131), the Chan Zuckerberg Initiative (2023-332162), the National Institutes of Health (NIH, U54NS093793), the Eunice Kennedy Shriver National Institute of Child Health and Human Development of the NIH (P50HD103555), the Chao Endowment, the Huffington Foundation, and the Jan and Dan Duncan Neurological Research Institute at Texas Children’s Hospital.
Author Competing Interests
The authors declare no competing interests.